# Stronger Than You Think: Benchmarking Weak Supervision on Realistic Tasks

### v1

@@ 3.4 Experimental Setup
We describe various aspects of our experimental setup: our pipeline, learning scenarios, a description of how we evaluate when WS is useful.
Pipeline. For backward compatibility, our experimental pipeline is based on WRENCH  [44]. This allows users to experiment with different LMs, EMs, and LFs in a standardized format, and with WS-specific tooling for dataset manipulation. We use the majority vote (MV) as the default LM for all of our WS experiments (following WRENCH’s takeaway, MV produce good if not the best performance on many of the benchmarking datasets). Additionally, we test the more complex COSINE [43] and ARS2 [39], which did not consistently improve over MV (see Appendix I).
For the end-model, we fine-tune a pretrained RoBERTa model in most cases [26] and also try BERT variants for domain-specific (i.e., LegalBERT[6] for Claude9 and Sci-BERT[4] for ChemProt) and non-English tasks (BERT-base-chinese, NB-BERT-base, and BERT-base-japanese [14, 29, 30] for MASSIVE18). Following the procedure used by Zhu et al. [47], we randomly select a set of hyperparameters from their provided hyperparameter search spaces. We finetune for a fixed 6,000 total steps and average each experiment over 5 trials. Notably, we can successfully replicate the results of [47] with our setup on existing WS benchmarks, as shown in Appendix E.
We run the experiments on NVIDIA A100, A40, A6000, A4000, and RTX-4090 GPUs.
Learning Scenarios. We include the three types of learning scenarios that we compare: supervised learning, WS learning, and CFT, all involving clean labels coming from only the validation set (as shown in Fig. 1).
  • Supervised: We use clean labels from the validation set directly for fine-tuning.
  • Weakly Supervised: We aggregate all weak labels from the training data with an LM, then fine-tune an EM using the training data with the aggregated labels. A standard WS pipeline uses data with clean labels from the validation set for hyperparameter searching or early stopping. This experiment varies the size of validation data for early stopping.
  • Continuous-Fine-Tuning: We further fine-tune the models trained using only weak supervision on the same set of clean validation labels.
Crossover Points. We solidify the concept of WS being useful by studying the ‘crossover points’ in performance over the validation set sizes, when comparing WS to other techniques. In our setting, crossover points we care about are between Supervised method and the WS/CFT. Crossover points, as an object of study, underpin our experimental results comparing WS to supervised learning.
@@ 4 Results and Analysis
4.1 Crossover Points
We define the crossover points in this set of plots as points where the supervised method (the green line) crosses with the CFT method (the blue line) in Figs. 4 and 3. This point shows the amount of data with clean labels needed for the supervised-only method to surpass the WS method that has access to the same labels as well as the LFs: the higher the crossover point, the more utility WS brings for that task.
We first conducted a crossover point analysis for the existing datasets from WRENCH [44] in Fig. 3, extending the results from [47]. We confirm that for most of these tasks, the crossover points are considerably smaller, less than 200 for 4 out of 6 tasks. Notably, these four datasets all have considerably smaller label cardinalities compared to the new datasets in BoxWRENCH (See Appendix D). These experiments verified the existing tasks are inadequate for benchmarking WS methods in realistic scnearios where WS holds significant utility over hand-labeling. We also performed the analysis on the named entity recognition datasets from WRENCH, with similar results (See Appendix H).
@@ 4.1 Crossover Points
Figure 3: Crossover points on six existing WS benchmarks: AGNews, Semeval, Yelp, ChemProt, Trec, IMDB. The crossover points for these tasks are low, but for four out of the six tasks, the crossover points are less than 200, which is substantially lower than the crossover points in BoxWRENCH datasets.
@@ 4.2 Improving LFs leads to higher crossovers
Previously, the LFs of ChemProt from Yu et al. [43], which are keywords-based, have a coverage of 0.86 on the test set, and a precision of 0.55 (i.e., the accuracy of on the set of examples where LM does not abstain).To improve the quality of the LF set, we made the following changes. First we changed keywords used in existing LFs. Next, we modified seven LFs by using the absolute difference in entity positions in the text as features. After making these changes, we had a coverage of 0.8102 and precision of 0.6321. When constructing these LFs, we consulted domain experts, (Jinoh Li, a Ph.D students in Neuroscience and Behavior at Harvard) that have sufficient Chemistry and Biology knowledge for insights on the LFs. The detailed modification is included in Fig. 13.
These modifications also resulted in higher accuracy for the continuously fine-tuned model given the same amount of labeled data. The accuracy crossover point is around 800 (in comparison to the old 600 as shown in Fig. 3), and the F1 score crossover point is around 1600, whereas the previous F1 score crossover points is around 800. ChemProt has high class imbalance, so the F1 score is an essential metric to consider.
This experiment demonstrates that more careful writing of LFs can yield higher crossover points. Current research benchmarks may suffer from suboptimal LF sets, and this example highlights how simple adjustments can significantly enhance results.
Figure 5: Crossover points on ChemProt. The new LF sets for ChemProt demonstrate a higher crossover point in both accuracy and F1-micro scores, reaching over 800 compared to the previous 600 as in Fig. 3. This shift suggests a notable improvement in performance.
4.3 Cross-lingual adaptability
For MASSIVE18, we consider the setting in which one has already trained a weakly supervised model based on the English version of the dataset and wishes to also be able to obtain a strong model for a different language. In doing so we consider the following approaches that leverage the portability of LFs which is a key advantage of WS deployments:
@@ 4.3 Cross-lingual adaptability
  • L2En-Inference-Supervised. This is an important baseline where we evaluate whether the English model alone can be directly used on other languages by simply first translating test examples to English. In Fig. 6, the validation size for this method represents the amount of clean data used to train the English model.
In this section, we present a set of experiments to demonstrate another dimension of WS’s usefulness: adaptability. Specifically, we look at this through the ability to re-use existing LFs in a multilingual setting. Here we present the accuracy of $f_{\theta_{\text{WS}}}$, $f_{\theta_{\text{CFT}}}$, and $f_{\theta_{\text{SUP}}}$ with our methods in Fig. 6, where different suffixes are used with each method for clarity. We use L2En-Inference-Supervised as a baseline, representing the use of non-language-specific models. For any given task, this method translate the data back to English and then use the existing English model. Fig. 6 demonstrates that language-specific models consistently achieve higher accuracy.
Figure 6: Crossover points on the multilingual MASSIVE18 dataset. The comparison of the green solid line and yellow solid line showed the importance of having a language specific model. Surprisingly using our method from Fig. 2(blue solid line) is able to achieve the same or even better performance compare to the light blue dotted line which used the weak labels from MASSIVE18.
The more realistic setting L2En-LFs Fig. 2 using back-translations also achieve comparable accuracy with Oracle-LFs which rely on the bijective relation of the original MASSIVE dataset, implying the usefulness of applying WS in real-world multi-language application. In addition, among all the datasets, we did not create new LFs in another languages; instead, we leveraged the generalizability of the existing LFs. As a result, the higher crossover points benefits from WS here are almost for free. We further noted that without developing language-specific models, the translation costs would be prohibitively high, because translation costs would incur at each inference time. We also noted that WS boost the crossover points to over 1000 even with the “free” noisy weak labels from our method.
Finally, we introduce a realistic setup in which we take 20% of the unlabeled training data from Norwegian, translate it to English, and apply the LFs. We then attach the weak labels generate by MV to the untranslated version. Additionally, we take 80% of the unlabeled training data with weak labels from English and translate them into Norwegian. These steps, combined with a Norwegian BERT model, leads to $f_{\theta_{\text{WS}}}$. The results is presented in Fig. 9, where the crossover points is above 1500, showcasing just by using existing LFs can provide incredible performance.
For languages where foundation models are less powerful, the techniques used by L2En-LFs can help to re-use the LFs and save a great amount of clean labels. This is helpful, especially when rarer languages have a limited clean data or even unlabeled data.
4.4 Label Model
@@
In our study, we conducted a comprehensive evaluation of WS using these LMs on our datasets. The performance of each LM was systematically assessed to determine its effectiveness in various scenarios. Detailed results of this evaluation are presented in Table 2, showcasing the comparative performance and highlighting the strengths of each model (See Appendix F for the full table with Amazon31 and Banking77).
For most of the new dataset, the CFT method outperforms the supervised-only method with a clear margin, especially in low-resource settings. Majority Vote and Dawid-Skene performed the best among the LM tested.
@@ 5 Limitations
There are several limitations to this work.
  • This work primarily focuses on text classification tasks, which are more common in practice; while other benchmarks such as WRENCH Zhang et al. [44] have both text classification and sequence-tagging tasks. Similar investigation and LF improvements could be confused on those tasks using our strategy and code base in further work.
  • This paper only primarily evaluated the performance of RoBERTa’s (some other BERT variant) EM. Similarly, this paper only tested the performance of some relatively simple LMs such as Majority Vote, Dawid-Skene, Snorkel, COSINE, and ARS2. More complex and recent models could be tested using the pipeline.
  • This paper fixed the number of steps and hyperparameter while training following settings from [47]. With proper early stopping and hyperparameter search, WS has the potential to achieve better results.
  • The MASSIVE18 dataset has one-to-one correspondence across languages, while in real-life scenarios, distribution shifts among different languages even on identical tasks. Thus multi-lingual dataset collected from real uses would be more effective in evaluating the performance
@@ 6 Conclusions
In this paper, we introduce BoxWRENCH, a benchmark that expands the evaluation of WS by addressing the limitations of existing benchmarks. By incorporating high-class cardinality, imbalance, and the need for domain expertise, BoxWRENCH  better reflects real-world data and tasks. Our results demonstrate that WS is highly effective, particularly in scenarios where traditional labeling is cost-prohibitive, and it achieves substantial performance gains before supervised methods surpass it at high label counts. We also establish that careful LF design and leveraging existing LFs in multilingual settings can significantly enhance WS applicability across diverse contexts. BoxWRENCH  sets a new standard for evaluating WS, with publicly released datasets, benchmarks, and tools to advance WS research and its practical deployment.
@@ Acknowledgements.
We are grateful for the support of the NSF under CCF2106707 (Program Synthesis for Weak Supervision) and the Wisconsin Alumni Research Foundation (WARF). Neel Guha is supported by the Stanford Interdisciplinary Graduate Fellowship.
@@ C.1 Original LFs
The original LFs have a coverage of 0.8637 and precision of 0.5512 on the covered data, with an accuracy of 0.4904 using Majority Vote and random tie-breaking (Wrench reported their accuracy in this way). These statistics are obtained from the test set.
@@ C.2 LF Improvement Details
Additionally, we developed a utility function, chemprot_enhanced, to extend the chemprot dataframe with two more columns: entity1_index and entity2_index. We improved our LFs to utilize these indices to check whether certain words occur between or near the two entities.
@@
Figure 7: Comparison of Chemprot results
@@ Appendix D Cardinality for WS tasks
+Majority vote 0.5652$\pm$0.0262 0.6665$\pm$0.0152 0.7519$\pm$0.0150 0.8218$\pm$0.0098 0.8654$\pm$0.0023
+Supervised Only 0.2362$\pm$0.0155 0.3777$\pm$0.0199 0.5705$\pm$0.0176 0.7522$\pm$0.0041 0.8488$\pm$0.0065
@@
+Majority vote 0.6850$\pm$0.0046 0.7103$\pm$0.0066 0.7429$\pm$0.0028 0.7706$\pm$0.0032 0.7923$\pm$0.0016
+Supervised Only 0.6603$\pm$0.0078 0.6999$\pm$0.0056 0.7357$\pm$0.0026 0.7673$\pm$0.0033 0.7902$\pm$0.0012
@@ Appendix H NER datasets
We included additional experiments related to named entity recognition datasets (BC5CDR[24], NCBI-Disease[11], CoNLL-03[36], OntoNotes 5.0[42]) at Fig. 11. Overall, we obeserve that crossover points are relatively small but note that each sequence has multiple label entities, which would further increase the cost of manual labeling compared to sequence classification tasks. Finding tasks and LFs for NER that have higher crossover points would be an interesting avenue for future work.

### v2

@@ 3.4 Experimental setup
We describe various aspects of our experimental setup: our pipeline, learning scenarios, and a description of how we evaluate when WS is useful.
Pipeline. For backward compatibility, our experimental pipeline is based on WRENCH  [46]. This allows users to experiment with different LMs, EMs, and LFs in a standardized format, and with WS-specific tooling for dataset manipulation. We use majority vote (MV) as the default LM for most of our WS experiments, following WRENCH’s takeaway that MV is one of the most consistent methods across various tasks. Appendix I includes results for COSINE [45] and ARS2 [41], but we did not find that they consistently improved upon MV. That being said, more effective WS methods in the future, would further strengthen our conclusions about WS’s effectiveness.
For the end-model, we fine-tune a pretrained RoBERTa model in most cases [27] and also try BERT variants for domain-specific (i.e., LegalBERT[6] for Claude9 and Sci-BERT[4] for ChemProt) and non-English tasks (BERT-base-chinese, NB-BERT-base, and BERT-base-japanese [14, 30, 31] for MASSIVE18). Following the procedure used by Zhu et al. [49], we randomly select a set of hyperparameters from their provided hyperparameter search spaces. Whenever training on weak labels, we do early stopping based on the validation data, following the practices of WRENCH, while we fine-tune for a fixed 6,000 total steps whenever training on clean labels, following Zhu et al. [49]. We run all the experiments on NVIDIA A100, A40, A6000, A4000, and RTX-4090 GPUs.
Learning Scenarios. We include the three types of learning scenarios that we compare: supervised learning, WS learning, and CFT, all involving clean labels coming from only the validation set (as shown in Fig. 1). Notably, we can successfully replicate the results of [49] with our setup on existing WS benchmarks, as shown in Appendix E.
  • Supervised: We use clean labels from the validation set directly for fine-tuning an end model.
  • Weakly Supervised: We aggregate all weak labels from the training data with an LM, then fine-tune an end model using the training data with the aggregated labels. Generally in WS, such as in WRENCH, the clean validation labels are used to guide hyperparameter search (for both the LM and end model). In our setup, we use it for early stopping.
  • Continuous-Fine-Tuning (CFT): We further fine-tune the models trained using only weak supervision on the same set of clean validation labels.
Crossover Points. To measure the usefulness of WS, we plot the performances of the three aforementioned methods for various amounts of clean validation data and inspect where their performance curves intersect. Specifically, the crossover point that we care about is between the supervised method, which uses only clean labels, and the better of WS and SFT, which make use of the weak labels.
@@ 4 Results and Analysis
4.1 Comparing crossover points between existing and new benchmarks
We first conducted a crossover point analysis for the existing datasets from WRENCH [46] in Fig. 3, extending the results from [49]. We confirm that for most of these tasks, the crossover points between Supervised (green) and CFT (blue) are quite small, less than 200 for four of six tasks. Notably, these four datasets all have considerably smaller label cardinalities compared to most datasets in BoxWRENCH (see Appendix D). We also performed the analysis on the named entity recognition datasets from WRENCH, with similar results (see Appendix H). Overall, these experiments verify that existing benchmarks are often inadequate for capturing realistic scenarios where WS holds significant utility over hand-labeling.
@@ 4.1 Comparing crossover points between existing and new benchmarks
Figure 3: Crossover points on six existing WS benchmarks: AGNews, Semeval, Yelp, ChemProt, Trec, IMDB. The crossover points for these tasks are low, less than 200 on four out of the six tasks, which is substantially lower than the crossover points in BoxWRENCH datasets.
[H]
@@ 4.2 Improving LFs leads to higher crossovers
Previously, the LFs of ChemProt from Yu et al. [45], which are keywords-based, have a coverage of 0.86 on the test set and a precision of 0.55 (i.e., the accuracy of the set of examples where LM does not abstain). To improve the quality of the LF set, we made the following changes. First, we changed the keywords used in existing LFs. Next, we modified seven LFs by using the absolute difference in entity positions in the text as features. After making these changes, the resulting LF set had a coverage of 0.81 and a precision of 0.63. When constructing these LFs, we consulted a Ph.D student in Neuroscience to leverage his domain expertise in Chemistry and Biology for insights on writing LFs. More details about the new LFs are included in Appendix C.
These modifications resulted in improved performance for CFT given the same amount of labeled data. When measuring F1, the crossover points grew to around 1600 as shown in Figure 5 (compared to 800 for the original LFs). When measuring accuracy, the improvement was smaller but consistent across the range of validation set sizes (see Fig. 7). This experiment demonstrates that more careful writing of LFs can yield higher crossover points. Current research benchmarks may suffer from suboptimal LF sets, and this example highlights how simple adjustments can enhance results.
Figure 5: Crossover points on ChemProt with our new LFs. The new LFs for ChemProt demonstrate a higher crossover point when measuring F1 performance.
4.3 Cross-lingual adaptability of LFs
In this subsection, we present a set of experiments to highlight an additional advantage of WS over manual labeling: its adaptability. Specifically, we examine the ability to reuse existing LFs in a multilingual setting on the MASSIVE18 dataset. We consider the setting in which one has already trained a weakly supervised model based on LFs written for the dataset’s English version and wishes also to obtain a strong model for other languages. In doing so, we consider the following approaches:
@@ 4.3 Cross-lingual adaptability of LFs
  • L2En-Inference-Supervised. This is an important baseline where we evaluate whether the English model alone can be directly used in other languages by simply first translating test examples into English. In Fig. 6, the validation size for this method represents the amount of clean data used to train the English model.
Here we present the accuracy of $f_{\theta_{\text{WS}}}$, $f_{\theta_{\text{CFT}}}$, and $f_{\theta_{\text{SUP}}}$ with our methods in Fig. 6, where different suffixes are used with each method for clarity. We use L2En-Inference-Supervised as an additional baseline, representing the use of non-language-specific models, which we observe underperforms training language-specific models (while also incurring additional costs at inference time).
Figure 6: Crossover points on the multilingual MASSIVE18 dataset. The comparison of the green and yellow solid lines showed the importance of having a language-specific model. Surprisingly, using our method from Fig. 2 (blue solid line) is able to achieve the same or even better performance compared to the light blue dotted line, which used the weak labels directly from the English version of MASSIVE18.
Meanwhile, we see that L2En-LFs demonstrates promise of adapting the English LFs, leading to crossover points of over 1000 in Chinese and Japanese, despite coming for “free” (i.e., without writing additional LFs in the target languages). It also achieves accuracy comparable to Oracle-LFs.
Finally, we also introduce a realistic setup in which even the amount of unlabeled data is limited for a rarer language, Norwegian. Specifically, we assume only a 20% subset of the Norwegian version of the MASSIVE is available, which we label using L2En-LFs. We then augment this data by translating the corresponding weakly labeled English examples into Norwegian for the remaining 80%. As shown in Fig. 9, this leads to a crossover points of above 1500 for fine-tuning Norwegian BERT (NB-BERT-base) [30]. This showcases that for rarer languages, for which foundation models may be less powerful and labels harder to obtain, adapting LFs can be of even larger value.
4.4 Label model ablations
@@
In our study, we conducted a comprehensive evaluation of WS using these LMs on our datasets. The performance of each LM was systematically assessed to determine its effectiveness in various scenarios. Detailed results of this evaluation are presented in Table 2, showcasing the comparative performance and highlighting the strengths of each model (see Appendix F for the full table with Amazon31 and Banking77). For most of the new datasets, the CFT method outperforms the supervised-only method with a clear margin, especially in low-resource settings. Majority Vote and Dawid-Skene performed the best among the label models tested.
@@ 5 Limitations
There are several limitations to this work. (1) We primarily focus on text classification tasks, which are more common in practice; while other benchmarks such as WRENCH Zhang et al. [46] include both text classification and sequence-tagging. Similar investigations and LF improvements would be valuable to study using our codebase in future work. (2) In our experiments, we used BERT-based end models and relatively simple label models such as Majority Vote, Dawid-Skene, Snorkel, COSINE, and ARS2. More recent end models and label models could also be worth testing with our pipeline. (3) We did not thoroughly tune hyperparameters while training weakly supervised models, following the setup from [49]. With more careful hyperparameter tuning, WS has the potential to achieve better results. (4) The MASSIVE dataset has one-to-one correspondences across languages, while in real-life scenarios, usage patterns and distribution shifts may exist across different languages, even on identical tasks.
@@ 6 Conclusions
In this paper, we introduce BoxWRENCH, a benchmark that expands the evaluation of WS by addressing the limitations of existing benchmarks. By incorporating high-class cardinality, imbalance, and the need for domain expertise, BoxWRENCH  better reflects real-world data and tasks. Our results show that on these more realistic tasks, weak supervision demonstrates significant utility, particularly in scenarios where traditional labeling is cost-prohibitive. We also show that careful LF design and adapting existing LFs in multilingual settings can significantly enhance the applicability of WS across diverse contexts. BoxWRENCH  sets a new standard for evaluating WS, with publicly released datasets, benchmarks, and tools to advance WS research and its practical deployment.
@@ Acknowledgements.
We are grateful for the support of the NSF under CCF2106707 (Program Synthesis for Weak Supervision) and the Wisconsin Alumni Research Foundation (WARF). Neel Guha is supported by the Stanford Interdisciplinary Graduate Fellowship. We thank Jinoh Lee for his contribution to LF improvements on ChemProt.
@@ C.1 Original LFs
The original LFs have a coverage of $0.864$ and precision of $0.551$ on the covered data, with an accuracy of 0.4904 using Majority Vote and random tie-breaking (WRENCH reported their accuracy in this way). These statistics are obtained from the test set.
@@ C.2 LF Improvement Details
Additionally, we developed a utility function, chemprot_enhanced, to extend the ChemProt dataframe in WRENCH format with two more columns: entity1_index and entity2_index. We improved our LFs to utilize these indices to check whether certain words occur between or near the two entities.
@@
Figure 7: This plot illustrates the crossover points when comparing ChemProt performance with our old and new LFs. The new LFs for ChemProt demonstrate a higher crossover point when measuring F1 performance and smaller (but consistent) gains in accuracy.
@@ Appendix D Cardinality for WS tasks
+Majority vote 0.565$\pm$0.026 0.667$\pm$0.015 0.752$\pm$0.015 0.822$\pm$0.010 0.865$\pm$0.002
+Supervised Only 0.236$\pm$0.016 0.378$\pm$0.020 0.571$\pm$0.018 0.752$\pm$0.004 0.849$\pm$0.007
@@
+Majority vote 0.685$\pm$0.005 0.710$\pm$0.007 0.743$\pm$0.003 0.771$\pm$0.003 0.792$\pm$0.002
+Supervised Only 0.660$\pm$0.008 0.700$\pm$0.006 0.736$\pm$0.003 0.767$\pm$0.003 0.790$\pm$0.001
@@ Appendix G MASSIVE60 and Using English for Augmentation
MASSIVE60, with our straightforward methodology, demonstrates a crossover point exceeding 500, highlighting its potential for the application of WS techniques.
@@ Appendix H NER datasets
We included additional experiments related to named entity recognition datasets (BC5CDR[25], NCBI-Disease[11], CoNLL-03[38], OntoNotes 5.0[44]) at Fig. 11. Overall, we observe that crossover points are relatively small but note that each sequence has multiple label entities, which would further increase the cost of manual labeling compared to sequence classification tasks. Finding tasks and LFs for NER that have higher crossover points would be an interesting avenue for future work.